OpenStudy (anonymous):

1/3(X + 9) = 12 27 1/3 9 12

OpenStudy (anonymous):

1/3(x + 9) = 12 -- distribute through the parenthesis 1/3x + 3 = 12 -- subtract 3 from both sides 1/3x = 12 - 3 1/3x = 9 -- divide by 1/3 x = 9/(1/3) x = 9 * 3/1 x = 27